# BPE: Behavioral Profiling Ensemble

### v1

@@ 3 BPE: Behavioral Profiling Ensemble
To address the limitations of conventional DES methods regarding storage overhead, retrieval latency, and failure in high-dimensional spaces, this section proposes BPE, a validation-free dynamic ensemble framework. The BPE framework abandons the traditional evaluation paradigm that relies on external validation sets for neighborhood retrieval. Instead, it adopts the concept of constructing model behavioral profiles via training set perturbations to reveal the intrinsic behavioral characteristics of base classifiers. Specifically, Section 3.1 elucidates the design philosophy of BPE, focusing on the paradigm shift from “resume screening” to “behavioral profiling.” Section 3.2 details a specific implementation algorithm, BPE-Entropy, based on information entropy. Finally, Section 3.3 provides a systematic analysis of the algorithmic complexity, demonstrating its advantages in computational efficiency and deployment flexibility.
@@ 3.1 Design Philosophy of BPE
The core design philosophy of the BPE framework lies in shifting the evaluation perspective of dynamic ensembles from external validation to internal consistency. Existing DES methods essentially follow a neighborhood-based validation paradigm, which assumes that the feature space surrounding a test sample shares similar decision boundaries with certain historical samples in the reference set. Under this assumption, the system infers a model’s performance on the current sample by retrieving neighbors from the reference set and calculating the model’s accuracy on these proxies.
However, this paradigm faces both theoretical and engineering challenges in practical applications. Theoretically, “distance” in high-dimensional feature spaces often loses its discriminative power [30], causing neighbors retrieved via Euclidean distance to be spurious (semantically irrelevant), thus introducing noisy evaluations. From an engineering perspective, reliance on a reference set necessitates maintaining an extensive historical database and performing expensive search computations during inference, which not only increases storage and latency overheads but also hinders the system’s ability to handle streaming or privacy-sensitive data.
To this end, we propose a novel paradigm centered on model behavioral profiling. We posit that the decision reliability of a model for a given instance should not be determined solely by its historical performance metrics, but rather by its intrinsic state when confronted with the current input. To capture these inherent behavioral attributes, we introduce training set perturbation as the core mechanism for establishing behavioral profiles. This approach is analogous to a “stress test” in psychology: rather than relying on a subject’s past records, we observe their authentic reactions under controlled environmental pressure.
Within the BPE framework, we construct a testing environment for profiling by injecting noise into the training set. Through this process, the divergence patterns in the model’s output distribution implicitly reveal its latent behavioral characteristics. Consequently, this allows us to discern when a model exhibits high uncertainty versus high predictive confidence. We allocate greater weight to a model when it demonstrates robust confidence. By calibrating a model’s response to different samples relative to its established behavioral profile, we derive a robust basis for dynamic weight assignment.
@@
    Minimal Storage Overhead: The maintenance cost is significantly lower than that of traditional DES methods, as BPE requires storing only a compact behavioral profile (typically a vector) for each model, rather than a raw reference dataset.
@@
    Inference Efficiency: Unlike retrieval-based DES methods, BPE necessitates only the calculation of lightweight statistical metrics during the inference phase. By eliminating complex nearest-neighbor search processes, BPE achieves a substantial advantage in inference speed, particularly in scenarios involving large-scale datasets.
Figure 1: Average Friedman ranks of BPE and twelve baseline models in homogeneous ensemble case.
@@ 3.2 BPE-Entropy
The use of negative entropy ensures a positive correlation between this metric and model confidence; that is, a larger $S$ value indicates a sharper predictive distribution and higher model certainty.
The first stage of the algorithm involves offline behavioral profiling, aimed at establishing a performance baseline for each model under simulated real-world conditions. Since profiling does not strictly rely on ground-truth labels, we move beyond traditional validation set retrieval and instead apply Gaussian perturbations to the training set feature space to simulate uncertainty in the testing environment.
@@
To implement this simulation while ensuring mathematical consistency, we explicitly handle nominal features prior to perturbation. Specifically, all categorical attributes are transformed into numerical vectors via One-Hot Encoding, ensuring the input $\mathbf{x}_{i}$ resides in a continuous metric space. The noise injection is applied to this encoded space. The selection of a Gaussian distribution for the noise vector $\epsilon$ is theoretically grounded in the Principle of Maximum Entropy proposed by Jaynes [31]. According to this principle, the Gaussian distribution has the maximum entropy among all distributions with a specified variance. Choosing this distribution ensures that we introduce the least amount of structural bias into the simulation, as it avoids implicit assumptions (such as the hard boundaries inherent in a uniform distribution) that are unjustified without prior knowledge. Consequently, for each sample $\mathbf{x}_{i}$, in the training set, we generate a perturbed sample $\tilde{\mathbf{x}}_{i}=\mathbf{x}_{i}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\delta^{2}\mathbf{I})$ represents the unbiased noise vector.
@@
where $\xi$ is a constant for numerical stability. The physical interpretation of $z_{k}$ is intuitive: it measures the number of standard deviations the current model confidence deviates from its historical behavioral profile. If $z_{k}>0$, the model exhibits higher-than-usual confidence and should be rewarded; otherwise, the model is considered confused and should be suppressed. Finally, we map these normalized scores to non-negative weights. To amplify the contribution of high-confidence models, we employ an exponential mapping function followed by normalization:
@@
where $\lambda$ is a sensitivity hyperparameter. The final ensemble output $H(\mathbf{x}_{\text{test}})$ is derived from the weighted combination of the base models. Owing to the versatility of the BPE framework, these weights can be applied directly to probability vectors or to ranking results, flexibly adapting to different task requirements:
@@
1:Training dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$; Pool of base classifiers $\mathcal{M}=\{h_{1},\dots,h_{K}\}$; Perturbation scale $\delta$; Sensitivity factor $\lambda$.
@@
16:  Compute unnormalized weight (Exponential Mapping):
17:   $w^{\prime}_{k}\leftarrow\exp(\lambda\cdot z_{k})$
18:end for
19:Normalize weights: $w_{k}\leftarrow w^{\prime}_{k}/\sum_{j=1}^{K}w^{\prime}_{j}$.
20:Return Ensemble output $H(\mathbf{x}_{\text{test}})=\sum_{k=1}^{K}w_{k}\cdot\Phi(\mathbf{x}_{\text{test}})_{k}$.
@@ 3.3 Complexity Analysis
To comprehensively evaluate the deployment potential of the BPE framework, this section analyzes its complexity in terms of space complexity and time complexity, contrasting it with conventional neighborhood-based DES approaches such as KNORA [8], LCA [7], and RRC [25].
Traditional DES methods necessitate maintaining the entire reference set (validation or training set) during the inference phase for retrieval purposes. Assuming a reference set size of $N$, a feature dimensionality of $D$, and $K$ base models, the space complexity of traditional DES is primarily dictated by the scale of the reference set, approximating $\mathcal{O}(N\cdot D+N\cdot K)$. In large-scale data scenarios, memory consumption grows linearly with sample accumulation, imposing significant constraints on storage-limited devices. In contrast, BPE compresses the distributional characteristics into concise “behavioral profiles” during the offline phase. For each base model, BPE solely requires storing two scalar statistics ($\mu_{k}$ and $\sigma_{k}$), resulting in an auxiliary storage overhead of merely $\mathcal{O}(K)$. Given that $K\ll N\cdot D$, BPE substantially lowers the barriers to deployment.
Regarding time complexity, it is essential to distinguish between offline profiling and online inference. In the offline phase, both traditional DES and BPE require a comprehensive inference pass over the dataset to compute statistical features. However, during the online inference phase, traditional DES must execute a $k$-nearest neighbor search for each test sample. Even with optimized indexing structures like KD-Trees, search efficiency in high-dimensional spaces degrades towards $\mathcal{O}(N\cdot D)$, causing inference latency to escalate significantly with the volume of historical data.
Conversely, BPE completely eliminates the dependence on historical data during inference, requiring only the computation of the current sample’s entropy and Z-score. The complexity of computing entropy is $\mathcal{O}(C)$ (where $C$ is the number of classes), and the complexity of normalization and weighting is $\mathcal{O}(K)$. Consequently, the total online selection complexity of BPE remains constant at $\mathcal{O}(K\cdot C)$. This implies that the inference speed of BPE is entirely independent of the scale of the training or reference data, maintaining constant computational overhead regardless of the magnitude of historical data.
Table 1: Complexity comparison between traditional KNN-based DES and the proposed BPE framework. Note that $N$, $D$, $K$, and $C$ denote the dataset size, feature dimension, number of models, and number of classes, respectively.
@@
Traditional DES High: $\mathcal{O}(N\cdot D)$ High: $\mathcal{O}(N\cdot K)$ High: $\mathcal{O}(N\cdot D)$ Linear
BPE (Ours) Low: $\mathcal{O}(K)$ High: $\mathcal{O}(N\cdot K)$ Fast: $\mathcal{O}(1)$ None (Constant)
4 Simulation Experiments
4.1 Experimental Setup
4.1.1 Data Generation
To systematically evaluate the performance of BPE across varying dimensionalities and data characteristics, we designed three synthetic experimental scenarios. All experiments utilize Monte Carlo simulation methods, with each scenario executing 1,000 randomized sampling iterations within specified parameter ranges to ensure the statistical robustness of the results.
Scenario 1: Binary Gaussian Mixture Classification. This scenario evaluates the algorithm’s performance in fundamental classification tasks. Data are generated using a multi-dimensional Gaussian Mixture Model (GMM). For each class $c\in\{0,1\}$, samples $\mathbf{x}$ are drawn from a mixture of $K$ Gaussian components:
$\mathbf{x}\sim\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})$ (7)
where $K$ represents the number of components, and $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Sigma}_{k}$ denote the mean vector and covariance matrix of the $k$-th component, respectively. To simulate observational noise inherent in real-world settings, we introduce label-flip noise into the training set. Let $y$ be the ground-truth label and $\tilde{y}$ be the observed label; the noise injection follows a Bernoulli distribution:
$P(\tilde{y}=y)=1-\rho,\quad P(\tilde{y}\neq y)=\rho$ (8)
In this experiment, the noise rate $\rho$ is sampled uniformly from the interval $[0.03,0.05]$, and the feature dimensionality $D$ is sampled uniformly from the integer set $\{10,\dots,30\}$.
Scenario 2: Multi-class Adaptability. This scenario examines the algorithm’s adaptability in complex multi-class environments. The data generation mechanism remains consistent with Scenario 1, with the exception that the number of classes $C$ is not fixed but follows a discrete uniform distribution $C\sim U\{3,10\}$. As $C$ increases, the overlapping regions in the feature space expand significantly, resulting in more intricate decision boundaries. This necessitates that the ensemble algorithm not only discriminates between positive and negative samples but also accurately quantifies the inter-class ambiguity via entropy.
Scenario 3: Non-linear Hastie Task. To verify the capability of the algorithm in handling non-linear manifolds, we employ the standard Hastie 10-2 dataset [30]. The decision boundary of this dataset is defined by a hypersphere and is linearly inseparable. For a 10-dimensional input vector $\mathbf{x}=(x_{1},\dots,x_{10})$, the label $y$ is determined by the following discriminant function:
$y=\begin{cases}1&\text{if }\sum_{j=1}^{10}x_{j}^{2}>\chi^{2}_{10}(0.5)\approx 9.34\\ 0&\text{otherwise}\end{cases}$ (9)
where $\chi^{2}_{10}(0.5)$ denotes the median of a chi-square distribution with 10 degrees of freedom. The concentric “onion-like” structure of this task presents a significant challenge to algorithms predicated on local linearity assumptions.
4.1.2 Base Model Pool

### v2

@@ 4 BPE: Behavioral Profiling Ensemble
To address the limitations of conventional DES/DCS methods regarding storage overhead, retrieval latency, and failure in high-dimensional spaces, this section proposes BPE, a validation-free dynamic ensemble framework. The BPE framework abandons the traditional evaluation paradigm that relies on external reference sets for neighborhood retrieval. Instead, it adopts the concept of constructing model behavioral profiles via training set perturbations to reveal the intrinsic behavioral characteristics of base classifiers. Specifically, Section 3.1 elucidates the design philosophy of BPE, focusing on the paradigm shift from “resume screening” to “behavioral profiling.” Section 3.2 details a specific implementation algorithm, BPE-Entropy, based on information entropy. Finally, Section 3.3 provides a systematic analysis of the algorithmic complexity, demonstrating its advantages in computational efficiency and deployment flexibility.
@@ 4.1 Design Philosophy of BPE
The core design philosophy of the BPE framework lies in shifting the evaluation perspective of dynamic ensembles from external validation to internal consistency. Existing DES/DCS methods essentially follow a neighborhood-based validation paradigm, which assumes that the feature space surrounding a test sample shares similar decision boundaries with certain historical samples in the reference set. Under this assumption, the system infers a model’s performance on the current sample by retrieving neighbors from the reference set and calculating the model’s accuracy on these proxies.
However, this paradigm faces both theoretical and engineering challenges in practical applications. Theoretically, “distance” in high-dimensional feature spaces often loses its discriminative power [32], causing neighbors retrieved via Euclidean distance to be spurious (semantically irrelevant), thus introducing noisy evaluations. From an engineering perspective, reliance on a reference set necessitates maintaining an extensive reference set and performing expensive search computations during inference, which not only increases storage and latency overheads but also hinders the system’s ability to handle streaming or privacy-sensitive data.
To this end, we propose a novel paradigm centered on model behavioral profiling. We posit that the decision reliability of a model for a given instance should not be determined solely by its historical performance metrics, but rather by its intrinsic state when confronted with the current input. To capture these inherent behavioral attributes, we introduce training set perturbation as the core mechanism for establishing behavioral profiles. This is analogous to a stress test: rather than relying on a subject’s past records, we observe their authentic reactions under controlled environmental pressure.
Within the BPE framework, we construct a testing environment for profiling by injecting noise into the training set. Through this process, the divergence patterns in the model’s output distribution implicitly reveal its latent behavioral characteristics. Consequently, this allows us to discern when a model exhibits high uncertainty versus high predictive confidence. We allocate greater weight to a model when it demonstrates consistent confidence. By calibrating a model’s response to different samples relative to its established behavioral profile, we derive a robust basis for dynamic weight assignment.
@@
    Minimal Storage Overhead: The maintenance cost is significantly lower than that of traditional DES/DCS methods, as BPE requires storing only a compact behavioral profile (typically a vector) for each model, rather than a raw reference dataset.
@@
    Inference Efficiency: Unlike retrieval-based DES/DCS methods, BPE necessitates only the calculation of lightweight statistical metrics during the inference phase. By eliminating complex nearest-neighbor search processes, BPE achieves a substantial advantage in inference speed, particularly in scenarios involving large-scale datasets.
Let $\mathcal{M}=\{h_{1},\dots,h_{K}\}$ denote a pool of probabilistic base classifiers. BPE is a framework consisting of three components: (i) a behavioral profile definition, which specifies the form of a model-specific profile $\mathcal{P}_{k}$ that summarizes the typical behavioral pattern of $h_{k}$; (ii) a profile construction rule, which defines how $\mathcal{P}_{k}$ is obtained from the model responses under a prescribed profiling field; and (iii) a deviation scoring rule, which computes a score to quantify how the behavior of $h_{k}$ on the current test instance deviates from $\mathcal{P}_{k}$. In Section 3.2 we present BPE-Entropy as one representative instantiation.
Figure 1: Comparison between BPE and traditional DES.
@@ 4.2 BPE-Entropy
Note that $S(\mathbf{p})$ is the negative entropy of $\mathbf{p}$, i.e., $S(\mathbf{p})=-H(\mathbf{p})$, so a larger $S$ value indicates a sharper predictive distribution and higher model certainty.
The first stage of the algorithm involves offline behavioral profiling, aimed at establishing a performance baseline for each model under simulated real-world conditions. Since profiling does not strictly rely on ground-truth labels, we move beyond traditional reference set retrieval and instead apply Gaussian perturbations to the training set feature space to simulate uncertainty in the testing environment.
@@
To implement this simulation while ensuring mathematical consistency, we explicitly handle nominal features prior to perturbation. Specifically, all categorical attributes are transformed into numerical vectors via One-Hot Encoding, ensuring the input $\mathbf{x}_{i}$ resides in a continuous metric space. The noise injection is applied to this encoded space. Although the perturbed one-hot vectors are no longer binary, this continuous relaxation is used only for probing model sensitivity during profiling and does not require semantic validity at the feature level. The selection of a Gaussian distribution for the noise vector $\epsilon$ is theoretically grounded in the Principle of Maximum Entropy proposed by Jaynes [33]. According to this principle, under a fixed variance constraint, the Gaussian distribution has the maximum entropy among all distributions. Choosing this distribution ensures that we introduce the least amount of structural bias into the simulation, as it avoids implicit assumptions (such as the hard boundaries inherent in a uniform distribution) that are unjustified without prior knowledge. Consequently, for each sample $\mathbf{x}_{i}$, in the training set, we generate a perturbed sample $\tilde{\mathbf{x}}_{i}=\mathbf{x}_{i}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\delta^{2}\mathbf{I})$ represents the unbiased noise vector.
@@
where $\xi$ is a constant for numerical stability (we set $\xi=10^{-12}$ in our experiments). The physical interpretation of $z_{k}$ is intuitive: it measures the number of standard deviations the current model confidence deviates from its historical behavioral profile. If $z_{k}>0$, the model exhibits higher-than-usual confidence and will be assigned a larger weight; otherwise, its contribution will be down-weighted. In practice, we clip $z_{k}$ to the range [5,5][-5,5] to avoid extreme values when $\sigma_{k}$ is small. Finally, we map these normalized scores to non-negative weights. To amplify the contribution of high-confidence models, we employ an exponential mapping function followed by normalization:
@@
where $\lambda$ is a sensitivity hyperparameter. Larger $\lambda$ yields a more peaked weight distribution. The final ensemble output $H(\mathbf{x}_{\text{test}})$ is derived from the weighted combination of the base models. Owing to the versatility of the BPE framework, these weights can be applied directly to probability vectors or to ranking results, flexibly adapting to different task requirements:
@@
1:Training dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$; Pool of base classifiers $\mathcal{M}=\{h_{1},\dots,h_{K}\}$; Perturbation scale $\delta$; Sensitivity factor $\lambda$; $\xi$ is a small constant $(10^{-12})$.
@@
16:  Clip: $z_{k}\leftarrow\min\big(\max(z_{k},-5),5\big)$.
17:  Compute unnormalized weight (Exponential Mapping):
18:   $w^{\prime}_{k}\leftarrow\exp(\lambda\cdot z_{k})$
19:end for
20:Normalize weights: $w_{k}\leftarrow w^{\prime}_{k}/\sum_{j=1}^{K}w^{\prime}_{j}$.
21:Return Ensemble output $H(\mathbf{x}_{\text{test}})=\sum_{k=1}^{K}w_{k}\cdot\Phi(\mathbf{x}_{\text{test}})_{k}$.
@@ 4.3 Complexity Analysis
To comprehensively evaluate the deployment potential of the BPE framework, this section analyzes its complexity in terms of space complexity and time complexity, contrasting it with conventional neighborhood-based DES/DCS approaches such as KNORA [10], LCA [9], and RRC [27].
Traditional DES/DCS methods necessitate maintaining the entire reference set during the inference phase for retrieval purposes. Assuming a reference set size of $N$, a feature dimensionality of $D$, and $K$ base models, the space complexity of traditional DES/DCS is primarily dictated by the scale of the reference set, approximating $\mathcal{O}(N\cdot D+N\cdot K)$. In large-scale data scenarios, memory consumption grows linearly with sample accumulation, imposing significant constraints on storage-limited devices. In contrast, BPE compresses the distributional characteristics into concise “behavioral profiles” during the offline phase. For each base model, BPE solely requires storing two scalar statistics ($\mu_{k}$ and $\sigma_{k}$), resulting in an auxiliary storage overhead of merely $\mathcal{O}(K)$. Given that $K\ll N\cdot D$, BPE substantially lowers the barriers to deployment.
Regarding time complexity, it is essential to distinguish between offline profiling and online inference. In the offline phase, both traditional DES/DCS and BPE require a comprehensive inference pass over the dataset to compute statistical features. However, during the online inference phase, traditional DES/DCS must execute a $k$-nearest neighbor search for each test sample. Even with optimized indexing structures such as KD-trees, the search cost often degrades to near-linear time in high-dimensional spaces, approaching $\mathcal{O}(N\cdot D)$ in practice, which can significantly increase inference latency as the reference set grows.
Conversely, BPE completely eliminates the dependence on historical data during inference, requiring only the computation of the current sample’s entropy and Z-score. The complexity of computing entropy is $\mathcal{O}(C)$ (where $C$ is the number of classes), and the complexity of normalization and weighting is $\mathcal{O}(K)$. Consequently, the total online selection complexity of BPE is $\mathcal{O}(K\cdot C)$. Importantly, this inference cost is independent of the scale of the training or reference data (i.e., independent of $N$), and thus remains constant with respect to dataset size.
Table 1: Complexity comparison between traditional KNN-based DES/DCS and the proposed BPE framework. Note that $N$, $D$, $K$, and $C$ denote the dataset size, feature dimension, number of models, and number of classes, respectively.
@@
DES/DCS High: $\mathcal{O}(N\cdot D)$ High: $\mathcal{O}(N\cdot K)$ High: $\mathcal{O}(N\cdot D)$ Linear
BPE Low: $\mathcal{O}(K)$ High: $\mathcal{O}(N\cdot K)$ Fast: $\mathcal{O}(K\cdot C)$ None
5 Experimental framework
5.1 Dataset
We selected 42 real-world datasets from OpenML for experimentation[34]. These datasets originate from diverse high-quality sources, including the UCI Machine Learning Repository and NASA, covering a broad spectrum of domains ranging from software engineering (software defect prediction) and healthcare (disease diagnosis and physiological monitoring) to natural science (high-energy physics, biological classification, and climate simulation), business finance (customer churn prediction and credit approval), as well as cybersecurity and information technology (spam and phishing website detection). Detailed specifications of these datasets are provided in Table 2. The configuration of the base model pool remains consistent with that of the simulation experiments. Furthermore, we conducted homogeneous model ensemble experiments; although such scenarios typically manifest as encapsulated algorithms in practical deployment (e.g., Random Forest), our homogeneous ensemble comprises 40 individual Decision Tree models, where the selection protocol for integration remains identical to that of the heterogeneous ensemble.
Table 2: Summary description of the 42 datasets used in the experiments.
Dataset ID Dataset Name #Ex. #Atts. #Num. #Nom. #Cl.
D1 9 autos 205 26 15 11 6
D2 10 lymph 148 19 3 16 4
D3 27 colic 368 23 7 16 2
D4 29 credit-approval 690 16 6 10 2
D5 30 page-blocks 5473 11 10 1 5
D6 42 soybean 683 36 0 36 19
D7 55 hepatitis 155 20 6 14 2
D8 59 ionosphere 351 35 34 1 2
D9 182 satimage 6430 37 36 1 6
D10 333 monks-problems-1 556 7 0 7 2
D11 337 SPECTF 349 45 44 1 2
D12 1017 arrhythmia 452 280 206 74 2
D13 1036 sylva_agnostic 14395 217 216 1 2
D14 1040 sylva_prior 14395 109 108 1 2
D15 1049 pc4 1458 38 37 1 2
D16 1050 pc3 1563 38 37 1 2
D17 1053 jm1 10885 22 21 1 2
D18 1063 kc2 522 22 21 1 2
D19 1067 kc1 2109 22 21 1 2
D20 1442 MegaWatt1 253 38 37 1 2
D21 1460 banana 5300 3 2 1 2
D22 1461 bank-marketing 45211 17 7 10 2
D23 1463 blogger 100 6 0 6 2
D24 1465 breast-tissue 106 10 9 1 6
D25 1467 climate-model-sim-crashes 540 21 20 1 2
D26 1480 ilpd 583 11 9 2 2
D27 1487 ozone-level-8hr 2534 73 72 1 2
D28 1489 phoneme 5404 6 5 1 2
D29 1494 qsar-biodeg 1055 42 41 1 2
D30 1496 ringnorm 7400 21 20 1 2
D31 1511 wholesale-customers 440 9 7 2 2
D32 4538 GesturePhaseSeg-Processed 9873 33 32 1 5
D33 6332 cylinder-bands 540 40 18 22 2
D34 40670 dna 3186 181 0 181 3
D35 40685 shuttle 58000 10 9 1 7
D36 40693 xd6 973 10 0 10 2
D37 40705 tokyo1 959 45 42 3 2
D38 40945 Titanic 1309 14 6 3 2
D39 40966 MiceProtein 1080 82 77 5 8
D40 40981 Australian 690 15 6 9 2
D41 41144 madeline 3140 260 259 1 2
D42 41981 SAT11-HAND-runtime-class 296 116 115 1 14
5.2 Base learner Pool